Calculus II, Spring 2022 (Thomas’ Calculus Early Transcendentals 13ed), http://www.math.nthu.edu.tw/˜ wangwc/

Brief solutions to selected problems in homework week 03

1. Section 10.5, problem 41:

Figure 1: Section 10.5, problem 41